OpenStudy (anonymous):

Solve the following system using any method. -3x + 4y = -2 12x - 16y = 8

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

no solution or ?

OpenStudy (superhelp101):

do you have any answer choices?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

-1, -2) (3,-2) (-3, 4) No solution. The system is inconsistent. Infinite number of solutions. The system is dependent.

OpenStudy (superhelp101):

yep I think no solution

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

12x-16y=8 -3x + 4y = -2 +3 +3 4y/4= -2+3x /4 y= -2/4 + 3x y= -0.5 +3x 12x- 16(-0.5 + 3x) = 8 12x + 8 -48x = 8 -36x + 8= 8 -8 -8 -36x = 0 No Solution

OpenStudy (anonymous):

& Just a little heads up, don't take someone's "guess" especially when your grade's on the line!